OpenStudy (anonymous):

i really need help i will give a medal Which is the equation of the line? A.2x – y = 3 B.x + 2y = –6 C.2x + y = –3 D.x – 2y = 6

OpenStudy (anonymous):

first of all, you rewrite all the equation on the above in this form y=mx+b, then it should be easy for you to find the slope and the y -int on the graph

OpenStudy (mathmale):

I'd suggest you start by finding the vertical and horiz. intercepts of the given graph. The horizontal intercept is ( what number ? , 0), and the vertical intercept is (0, what number?)

OpenStudy (anonymous):

find the slope and the y-int on the graph first

OpenStudy (mathmale):

How colorful, chloe! Two different approaches, both equally valid. Take your pick!

OpenStudy (anonymous):

the thing is when ever i move to one part of this i completely forget the next >.<

OpenStudy (mathmale):

I promise, it gets easier as you continue to practice.

OpenStudy (anonymous):

um so the slope would be -3/6?

OpenStudy (anonymous):

yap, simplify the slope,

OpenStudy (anonymous):

that would be -1/2

OpenStudy (anonymous):

or -.5 (-.50)

OpenStudy (anonymous):

hold on, the slope should be run downward -3 and leftward -6

OpenStudy (anonymous):

no, you don't decimal the slope

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

did you find the slope? now try to find the y-int,

OpenStudy (mathmale):

@colorfulchloe : Please note that the graph rises as we move from left to right. In other words, the slope of this line is positive. Two points on the graph that are easy to read are (6,0) and (0,-3) You could apply the slope formula here. Note that your result must be a positive number in this example.

OpenStudy (anonymous):

Oooh! The Y intercept is -3 and the x is 6

OpenStudy (anonymous):

and then the whole thing put together would be -3=(-1/2)6+-3 ?

OpenStudy (anonymous):

good job! now, you got slope and the y-int, try to plug in the number in this form y=mx+b, you dont plug in the x-int, just the slope=m and the y-int=b

OpenStudy (anonymous):

hold on, look at your slope again, it should be positive, as i had explained before that on the graph, the slope is going down 3 and left 6, so what is the slope ?

OpenStudy (anonymous):

oh so it would just be 1/2 and not -1/2

OpenStudy (anonymous):

yap, you got it

OpenStudy (anonymous):

so write out the hold thing in this form y=mx+b, just plug in the slope and y-int.

OpenStudy (anonymous):

that would be -3=1/2*6+(-3) right?

OpenStudy (anonymous):

no, you have to write in an equation form, y=mx+b, just plug in the number slope and the y-int. in the question you dont' need the x-int, just ignore that.

OpenStudy (anonymous):

so leave where the 6 is "x"

OpenStudy (anonymous):

and make it -3=1/2*x+(-3)

OpenStudy (mathmale):

You've found that the slope, m, is 1/2, and that the y-intercept, b, is (0,-3). Merely substitute these values into the equation y = m*x + b.

OpenStudy (anonymous):

and then leave 6 out -x intercept-

OpenStudy (anonymous):

y=(slope)*x+(y-int)

OpenStudy (anonymous):

and then i solve

OpenStudy (anonymous):

the equation form of this graph should be y=1/2x-3

OpenStudy (mathmale):

Actually, you "solve" for the slope, m, and the y-intercept, b, and then SUBSTITUTE those values into y = m*x + b. Try it here and now, please.

OpenStudy (anonymous):

ok i got it sorry to use up your time, both of you.